OpenStudy (anonymous):

PLEASE HELP The current in a nearby river is 5 mph. John wants to be able to travel up the river for a distance of 30 miles and return to his starting point in a total of 8 hours. How fast must John be able to row his boat in order to accomplish this? (Assume that the speed that John can row refers to how fast he can row in still water.)

OpenStudy (anonymous):

@dan815 @nincompoop

OpenStudy (mathstudent55):

He travels a total time of 8 hours. If he travels up the river for t time, how much time will he travel down the river?

OpenStudy (anonymous):

i have no idea!! please explain

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

Ok. The total trip is 8 hours. time going up + time going down = 8 hours If you let the time going up be called just t, then you have t + time going down = 8 hours Ok so far?

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

Now let's see if we can get an expression for the time going down. t + time going down = 8 Subtract t from both sides: time going down = 8 - t Now we have: Time going up = t Time going down = 8 - t Ok?

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

Ok. I put the info we have in the table below. Now we need the speed going up and the speed going down, and the distances going up and down. 

OpenStudy (mathstudent55):

Let's do the distances first bec they are very easy.

OpenStudy (mathstudent55):

How far does he row each way?

OpenStudy (anonymous):

like how am i supposed to know im so confused on this question

OpenStudy (mathstudent55):

Have you read the problem? It's written right there. There is no calculation needed.

OpenStudy (mathstudent55):

Does this help? \(\sf John ~wants ~to ~be ~able ~to ~travel ~up ~the ~river ~for ~a ~distance ~of ~\huge \color{red}{30 ~miles}\) \( \sf and ~return ~to ~his ~starting ~point ~in ~a ~total ~of ~8 ~hours. \)

OpenStudy (anonymous):

so for up its 30 ?

OpenStudy (mathstudent55):

Right, Each part of the trip is 30 miles. We can add the distances to our table.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):